OpenStudy (anonymous):

The three sides of a triangle are 17, 23 and 30. Correct to the nearest tenth of a degree, what is the measure of the angle opposite to the largest side?

OpenStudy (wolf1728):

hgt² = 17² - x² A) hgt² = 289 - x² hgt² = 23² -y² B) hgt² = 529 -y² We can see from the diagram that x = 30 -y so we can input this into equation A hgt² = 289 - (30 -y)² hgt² = 289 -900 +60y -y² both equation A and B are both = hgt² so we can make these equal to each other 529 -y² = 289 -900 +60y -y² 1140 = 60y y = 19 and therefore x = 11 arc sine (a) = 11/17 angle a = 40.32021507 arc sine (b) = 19/23 angle b = 55.6988472432 angle opposite the side of length 30 = 40.32 plus 55.70 = 96.0 degrees